A ball is thrown forward from the top of a cliff with a velocity of 10ms. The height of the cliff above the ground is 45m. Calculate the distance of the ball to the hirizontal just before hitting the ground Please answer

Grade:12

3 Answers

Saurabh Koranglekar
3 years ago
Dear student

Since the velocity is horizontal

Vertical velocity can be assumed to be 0

Time required to reach ground is (2*45/g)^0.5 = 3 sec

Now horizontal dist = 10*3 = 30 meters

Regards
Sandeep
3 years ago
The answer would be 30 m as ....
the time taken by the ball toreach the drowns is 3 second by half at square =45
so 3*10= 30............................ Answer